Linux Meta-FAQ

Version 3.07 June 11, 1994

This is the Meta-FAQ for Linux. It is mainly a list of valuable sources of information. Check these sources out if you want to learn more about Linux, or have problems and need help. Lars Wirzenius (wirzeniu@cc.helsinki.fi) wrote the first version of this document, and it is now maintained by Michael K. Johnson (johnsonm@Sunsite.unc.edu). Mail me if you have any questions about this document.

NOTE: Filenames in this article are for the tsx-11.mit.edu ftp site unless otherwise noted. (see below for names of more ftp sites). Files are usually located in similar places on other sites. The names are relative to the directory /pub/linux/ on tsx-11.

What is Linux? Linux is an independent implementation of the POSIX operating system specification, with SYSV and BSD extensions, that has been written entirely from scratch. It has no proprietary code in it. Linux is freely distributable under the GNU Public License. It only works on IBM PC compatibles with an ISA or EISA bus and a 386 or higher processor. See the FAQ for more exact hardware requirements. The Linux kernel is written by Linus Torvalds (torvalds@kruuna.helsinki.fi) from Finland, and by other volunteers. Most of the programs running under Linux are generic Unix freeware, many of them from the GNU project.

The Linux FAQ A collection of common problems and their solutions. Answers many questions faster than the net. Stored on many Linux ftp sites (docs/) and rtfm.mit.edu, the general archive site for all FAQs.

The Linux HOWTO’s These are somewhat like FAQ’s, but instead of answering common questions, they explain how to do common tasks, like ordering a release of Linux, setting up print services under Linux, setting up a basic UUCP feed, etc. See sunsite.unc.edu, directory /pub/Linux/docs/HOWTO/ for all the HOWTO’s.

Linux newsgroups There are several Usenet newsgroups for Linux. It is a good idea to follow at least comp.os.linux.announce if you use Linux. Comp.os.linux.announce is moderated by Matt Welsh and Lars Wirzenius. To make submissions to the newsgroup, send mail to linux-announce@tc.cornell.edu. You may direct questions about comp.os.linux.announce to Matt Welsh, mdw@sunsite.unc.edu The newsgroup comp.os.linux.admin is an unmoderated newsgroup for discussion of administration of Linux systems. The newsgroup comp.os.linux.development is an unmoderated newsgroup specifically for discussion of Linux kernel development. The only application development questions that should be discussed here are those that are intimately associated with the kernel.

The newsgroup comp.os.linux.help is an unmoderated newsgroup for any Linux questions that don’t belong anywhere else.

The newsgroup comp.os.linux.misc is the replacement for comp.os.linux, and is meant for any discussion that doesn’t belong elsewhere.

In general, do not crosspost between the Linux newsgroups. The only crossposting that is appropriate is an occasional posting between one unmoderated group and comp.os.linux.announce. The whole point of splitting comp.os.linux into many groups is to reduce traffic in each. Those that do not follow this rule will be flamed without mercy…

Other newsgroups Do not assume that all your questions are appropriate for a Linux newsgroup just because you are running Linux. Is your question really about shell programming under any unix or unix clone? Then ask in comp.unix.shell. Is it about GNU Emacs? Then try asking in gnu.emacs. Also, if you don’t know another group to ask in, but think there might be, politely ask in your post if there is another group that would be more appropriate for your question. At least the groups comp.unix.{questions,shell,programming,bsd,admin}, and comp.windows.x.i386unix should be useful for a Linux user.

The World-Wide Web Matt Welsh, mdw@sunsite.unc.edu, maintains the home WWW page for the Linux project. The URL is http://sunsite.unc.edu/mdw/linux.html

The Linux Journal A new magazine called Linux Journal was recently launched. It includes articles intended for almost all skill levels, and is intended to be helpful to all Linux users. Subscriptions are 19 US Dollars worldwide. Subscription inquiries can be sent via email to subs@ssc.com or faxed to (U.S.) 1-206-527-3385 or mailed to Linux Journal, PO Box 85867, Seattle, WA 98145-1867 USA. Please do not send credit card numbers via email; the internet is not secure, and it is entirely possible that a technologically adept thief may steal your credit card number and cost you a large sum of money if you do.

1 Getting Linux

1.1 Linux FTP sites

A more complete list of Linux FTP sites is in the Linux INFO-SHEET (docs/INFO-SHEET). The most important sites are listed here; please see the INFO-SHEET for a site nearer to you (there are many mirrors).

Textual name Numeric address Linux directory
tsx-11.mit.edu 18.172.1.2  /pub/linux
sunsite.unc.edu 152.2.22.81  /pub/Linux
nic.funet.fi  128.214.6.100  /pub/OS/Linux

These sites are the main “home” sites for Linux where most uploads take place. There are many mirror sites; please use the closest (network-wise) site to you.

1.2 Linux on BBS’s

Occasionally, someone posts a list of BBS’s that have Linux available for download. Try them if you can’t FTP. This list is available by ftp as docs/bbs.list at tsx-11. You may use ftpmail (described elsewhere in this document) to get this list if you have mail access, or ask a friend with internet access to get it for you.

1.3 Linux on physical media

Linux is distributed on physical media, including floppies, CD-ROM, and tape, by several commercial vendors. Please read the distribution HOWTO, posted regularily to comp.os.linux.announce, and available for ftp at sunsite.unc.edu as /pub/Linux/docs/HOWTO/distribution-HOWTO.

Also check advertisements/* at tsx-11 for advertisements of other ways to get linux, including a list of people that will make copies for you at low prices. Also available there are advertisements for other commercial services related to Linux.

1.4 Commercial networks

GEnie mirrors most of tsx-11 and sunsite. CompuServe has only very limited very old Linux archives.

1.5 Mailservers and such

Sunsite offers ftp-mail service — mail ftpmail@sunsite.unc.edu.

2 Linux distributions (aka “releases”)

Linux is distributed by its author only as a kernel. Other people have put together “distributions” that can be used. These distributions pair the kernel as released by the author with software, to make a complete working package. Most releases include application programs as well as system software, providing “one stop shopping” for Linux.

H.J. Lu (the Linux GCC maintainer) also has a set of disks, including a combined boot and rootdisk, and some additional disks with more programs. This package assumes you are already familiar with Linux. See directories packages/GCC/rootdisk and packages/GCC/basedisk on tsx-11.

There are several distributions of Linux, which are available at various sites. Sunsite mirrors almost all the distributions in /pub/Linux/distributions. The recommended distributions are Slackware, MCC, and soon, Debian. These are all available for free over the internet, and some are also sold on floppies and CD-ROM’s.

3 Linux mailing-lists

Used mostly for discussion between developers of new features and testers of pre-release versions. See addresses in the FAQ. Many lists are based in the Mail-net system at Helsinki University of Technology. Send mail to linuxactivists-request@niksula.hut.fi, and you will get mail explaining how to get any of the many “channels” available. Save this mail, as it tells you how to unsubscribe to the channels, and if you post annoying messages to the rest of the list complaining about not being able to get off the list (because you didn’t follow instructions and save the mail telling you how to unsubscribe), you will likely be flamed for wasting international bandwidth and money.

4 Documentation for various programs

Many programs come with some sort of documentation, often in a file called README or something similar. It is a VERY good idea to read them with care. It is boring to see (and answer) questions that are answered in the documentation.

5 More Documentation

The Linux Documentation Project is working on a lot of documentation. Already, over 600 pages of book-style documentation has been released to the general public, and a large set of man pages has also been released, with more to follow. Check sunsite.unc.edu:/pub/Linux/docs/LDP for documents written by the LDP.

6 Keeping track of current releases

New releases, programs, and ports are usually announced in comp.os.linux.announce. Finger torvalds@klaava.helsinki.fi to get some information about the current kernel (often long!).

7 The Linux Hardware Compatibility List

Posted occasionally to comp.os.linux.announce, and found on the major Linux FTP sites (tsx-11:docs/compat.list). This list lists some of the devices and hardware that Linux currently supports. It’s useful for anyone wanting to purchase or upgrade their system. Be aware that it cannot be complete, as there is more hardware out there than Linux users can own.

8 Legalese

Trademarks are owned by their owners. Satisfaction not guaranteed. No warranties about this document. Void where prohibited.